OpenStudy (anonymous):

A rock salt (NaCl), ice, and water mixture is used to cool milk and cream to make homemade ice cream. How many grams of rock salt must be added to water to lower the freezing point 10.7°C?

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{\theta}\) where, \(\theta=b\) when finding BP, and \(\theta=f\) when finding freezing point (By that i mean use the appropriate constant.) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362